# A Review of Online Diffusion Policy RL Algorithms for Scalable Robotic Control

### v1

@@ 4.1.3 Experimental Design
Environmental Parallelization Capability The efficiency of online learning is often dictated by the degree of environment parallelization, which can be limited by the availability of computational resources or the complexity of the dynamics itself, such as large-scale diffusion architectures [rdt1b, diffusionvla, hybridvla, dexvla] or the complexity of exteroceptive observations (e.g., high-resolution RGB images).
@@
PPO [ppo] DDPG [ddpg] TD3 [td3] SAC [sac] DIPO [dipoerror] QVPO [qvpoerror] DACER [dacer] GenPO [genpo]
@@
4.2 Experiment 1: General Performance w.r.t. Various Tasks
@@
In environments with extensive agent parallelization, proximity-based on-policy algorithms, specifically PPO and GenPO, demonstrate superior overall performance. While PPO serves as a robust and high-performance baseline, GenPO exhibits a higher performance ceiling, securing the top rank in 6 out of 12 tasks (compared to 5 for PPO). Although GenPO showed lower performance than PPO in specific tasks such as Humanoid, Allegro Cube, and G1 Rough, it achieved significant performance gains—ranging from $1.51\times$ to $5.43\times$—in tasks like Franka Lift, Anymal-D, and Unitree Go2. These results underscore the potential of on-policy DPRL to transcend the representational limitations of conventional Gaussian policies in complex robotic tasks.
Conversely, off-policy algorithms generally exhibit slower convergence and suboptimal asymptotic performance compared to their on-policy counterparts. This performance gap is primarily attributed to the challenge of tracking rapidly shifting data distributions within replay buffers under high-throughput parallelization [genpo]. However, we observe a notable advantage of incorporating diffusion policies in complex locomotion tasks. Among the off-policy DPRL methods, DIPO achieved a mean rank (MR) of 3.58, significantly outperforming traditional off-policy algorithms such as DDPG (MR=6.58), TD3 (MR=5.50), and even SAC (MR=4.33), which is widely considered a strong classical baseline in off-policy RL. This performance gap demonstrates that integrating Diffusion models into the reinforcement learning framework yields substantial and meaningful performance gains. This suggests that the expressive power of Diffusion Models provides a more resilient policy representation for high-dimensional, complex robotic control.
@@
DIPO [dipoerror] QVPO [qvpoerror] DACER [dacer] GenPO [genpo]
@@
A closer examination of the four DPRL paradigms reveals distinct limitations: DIPO consistently outperforms QVPO and DACER. The suboptimal performance of QVPO (MR=5.33) is likely due to its sensitivity to reward scaling; since the $Q$-related reweighting function is directly coupled with the magnitude of manually designed rewards, it struggles with inconsistent reward landscapes across different tasks. DACER recorded the lowest performance (MR=6.67), failing to converge in the majority of tasks. This failure can be attributed to the inherent high variance of policy gradients [pgvar] and the extreme difficulty of optimizing diffusion models directly through BPTT using noisy Q-values as a supervisory signal, which often leads to unstable gradient dynamics [bengio1994learning, ddpograd].
4.3 Experiment 2: Parallelization Capability
These findings suggest a clear trade-off: while GenPO and PPO are optimized for high-throughput scalability, DIPO is uniquely suited for scenarios where extensive parallelization is computationally prohibitive. This is particularly relevant for training overparameterized models, such as Vision-Language-Action (VLA) architectures [openvla], where the memory and computational costs of maintaining thousands of concurrent environments are often infeasible.
Furthermore, our analysis highlights a critical gap in current online DPRL literature, which predominantly reports performance at a single, fixed scale [qsm, dipoerror, qvpoerror, dpmdsdac, dacer]. Our results underscore that environment scale is not merely a hyperparameter but a decisive factor in algorithmic ranking. Consequently, future benchmarking in the DPRL domain should incorporate diverse parallelization scales to ensure a comprehensive assessment of both scalability and robustness.
@@ 4.3 Experiment 2: Parallelization Capability
PPO [ppo] 874.32 1.201 0.555 470.40 1.031 0.517 4.92 1.630 1.287
DDPG [ddpg] 629.96 1.580 1.806 539.40 1.607 2.026 96.04 1.619 1.599
TD3 [td3] 580.63 1.595 1.859 559.83 1.635 2.002 90.61 1.652 1.498
SAC [sac] 610.05 1.552 1.787 583.63 1.504 1.960 99.49 1.649 1.645
DIPO [dipoerror] 546.78 1.518 1.816 530.75 1.641 2.016 97.49 1.548 1.617
QVPO [qvpoerror] 564.87 1.597 1.820 564.27 1.545 2.045 98.62 1.610 1.581
DACER [dacer] 644.37 1.569 1.825 552.73 1.587 2.023 88.93 1.530 1.562
GenPO [genpo] 961.39 0.495 0.344 862.80 0.459 0.400 58.30 0.786 0.547
@@
PPO [ppo] 423.68 1.377 0.776 789.82 1.330 0.600 858.58 1.357 0.640
DDPG [ddpg] 64.06 1.627 1.996 690.03 1.534 1.814 651.14 1.478 1.821
TD3 [td3] 70.54 1.670 1.880 731.15 1.522 1.797 582.41 1.543 1.770
SAC [sac] 89.21 1.571 1.860 671.47 1.622 1.815 518.65 1.511 1.869
DIPO [dipoerror] 83.36 1.719 1.917 685.57 1.599 1.788 474.08 1.604 1.845
QVPO [qvpoerror] 86.36 1.580 1.937 704.72 1.568 1.828 510.62 1.490 1.855
DACER [dacer] 58.93 1.660 2.000 719.18 1.525 1.799 625.11 1.567 1.798
GenPO [genpo] 869.12 0.596 0.431 568.34 0.850 0.568 493.95 0.992 0.521
@@
4.4 Experiment 3: Diffusion Step Scalability
@@
Conversely, DACER and GenPO, which perform backpropagation through all timesteps during training, exhibit a distinct inverse scalability pattern. While GenPO achieves state-of-the-art performance at lower step counts (KK==5, 10), its performance deteriorates significantly as $K$ increases, eventually suffering a catastrophic drop at KK==50. We attribute this failure to the numerical instability and gradient vanishing inherent in backpropagating through an excessively long denoising chain [bengio1994learning, gradclip]. Indeed, training diffusion models via BPTT is notoriously difficult, as the recursive nature of the denoising process compounds gradient-related issues as the chain length grows [ddpograd]. Interestingly, DACER exhibits the most erratic behavior; despite being a BPTT-based method, it fails to achieve meaningful reward levels regardless of the $K$ value. This suggests that its BPTT objective may fundamentally have a scalability issue with end-to-end diffusion optimization in online DPRL tasks.
4.5 Experiment 4: Cross-embodiment Generalizability
@@
4.6 Experiment 5: OOD Environment Generalizability
@@ 5 Discussion
5.1 Topic 1. What are the current challenges of the online DPRL?
Our empirical analysis in Sec. 4 reveals several fundamental bottlenecks inherent in existing online DPRL paradigms. First, DIPO, one of the representative algorithms of action gradient methods, often suffers from suboptimal performance due to the difficulty of obtaining accurate gradients from a potentially biased Q-critic [dppo, qvpoerror]. Specifically, updating actions via deterministic gradients and re-storing them into the replay buffer in DIPO may violate the Markov Decision Process (MDP) assumptions, as the stored transitions no longer reflect the current environment-agent interaction distribution [DDiffPG].
Second, QVPO, categorized by $Q$-weighting methods, exhibits vulnerability to reward scale sensitivities, as the $Q$-related reweighting function is directly coupled with the magnitude of manually designed rewards. For example, the derived Q-weighted variational objective for diffusion policy training cannot handle negative rewards properly [dpmdsdac].
Third, while the Proximity-based method, GenPO, benefits from large-scale parallelization in simulation, its performance degrades significantly in resource-constrained scenarios (refer to Sec. 4.3). Moreover, the intractability of the exact log-likelihood in diffusion models necessitates crude approximations or specialized architectures like invertible diffusion [edict], which imposes a bottleneck on both representational flexibility and inference speed.
Lastly, as demonstrated in Sec. 4.4, BPTT-based methods such as GenPO and DACER encounter severe scalability issues when the number of diffusion steps is increased. This is primarily due to gradient instability—including vanishing and exploding gradient—which inherently arises when backpropagating through an extensively long denoising chain [bengio1994learning, gradclip]
In summary, while each online DPRL paradigm offers unique theoretical advantages, they collectively face significant hurdles—ranging from gradient inaccuracy and reward-scale sensitivity to computational intractability and backpropagation-related scalability limits. Addressing these intertwined challenges is essential for transitioning from specialized simulation environments to reliable, high-performance robotic control in diverse real-world scenarios. Moreover, as simulation environments become increasingly efficient at supporting large-scale parallelization [isaaclab, brax], the demand for robust on-policy algorithms is expected to grow. Given the current scarcity of literature on on-policy DPRL compared to its off-policy counterparts, we encourage future researchers to explore this relatively uncharted territory to unlock the full potential of high-throughput robotic learning.
5.2 Topic 2. How do we select the online DPRL algorithms in specific robotic scenarios?
@@
5.3 Topic 3. What is the Future Avenue of Online DPRL?
@@
While several studies [dppo, parl, pgd, polygrad] have explored integrating online DPRL with offline-to-online and model-based approaches, the field remains in its early stages with numerous open challenges. To advance online DPRL toward more general robotic applications, we identify several promising research directions that leverage established RL paradigms, offering guidance for future investigation.
5.3.1 Action Chunking and Trajectory Planning
@@
While existing online DPRL methods focus on single-action per-predictions (as policies), real-world robotics often employs planning-based approaches that generate action chunks [actionchunking] or multi-step trajectories [openvla] (as planners). Recent work on Q-Chunking [qchunking] has demonstrated Q-learning extensions to action sequences for offline-to-online RL. Integrating temporal action chunking into online model-free DPRL could bridge reactive control and deliberative planning by leveraging diffusion models’ multi-step prediction capabilities to generate temporally consistent action sequences aligned with RL objectives. This evolution from state-to-action policies to state-to-trajectory planners would enhance sample efficiency and robustness in long-horizon tasks.
5.3.2 Safe Reinforcement Learning Integration
@@
Our experimental analysis (Sec 4.6) shows that simulation-trained policies can exhibit overly aggressive behaviors from overfitting to source domains when maximizing rewards without safety constraints. Safe RL approaches—including constraint-based methods (e.g., Constrained Policy Optimization [cpo]), barrier functions [cbf], and robust control [robustrl]—could naturally integrate with online DPRL to provide formal safety guarantees during exploration while maintaining diffusion-based action expressiveness, facilitating safer deployment in safety-critical applications.
5.3.3 Multi-Agent Reinforcement Learning
@@
Multi-agent scenarios present exponentially greater complexity than single-agent settings due to non-stationarity, credit assignment, and coordination requirements. However, advances in large-scale physics simulators capable of efficiently simulating thousands of agents make online Multi-Agent DPRL (MADPRL) particularly promising. Initial approaches could integrate established MARL algorithms [mappo, maddpg, qmix] with online DPRL frameworks, while sophisticated extensions might leverage diffusion models’ ability to capture complex joint action distributions for coordination strategies or employ centralized training with decentralized execution paradigms.
5.3.4 Inverse Reinforcement Learning from Demonstrations
@@
Inverse Reinforcement Learning (IRL) [irl, maxentirl, gail] infers reward functions from expert demonstrations, solving the inverse problem of standard RL. Since most diffusion policies [diffusionpolicy, mpd, diffuser] research relies on imitation learning from demonstrations, a natural extension reformulates these as reward learning problems rather than direct behavioral cloning. This “demonstration-to-reward-to-policy” paradigm offers key advantages: learned rewards generalize beyond demonstrated state-action pairs, provide clear objectives for online fine-tuning, and extract value from suboptimal demonstrations. Integrating IRL with online DPRL enables a refined learning approach where offline behavioral priors are optimized through interaction, potentially surpassing both standard imitation and pure offline-to-online approaches.
5.3.5 Hierarchical Reinforcement Learning
@@
Hierarchical RL (HRL) [hrl, hiro, latenthrl] decomposes complex decisions into multi-level structures where higher-level policies select among lower-level sub-policies to accomplish long-horizon tasks. This offers compelling synergy with diffusion policies: while diffusion models excel at generating diverse, multimodal actions, their stochastic nature introduces substantial variance. A hierarchical framework could employ a simpler higher-level policy to select among trajectory samples from lower-level diffusion sub-policies, mirroring batch action sampling [diffusionql, qvpoerror] but within a principled HRL structure. This transforms the exploration-exploitation trade-off from action space to a more manageable sub-policy selection problem, enabling sample-efficient learning of compositional behaviors while preserving diffusion models’ expressive capacity.
6 Conclusion

### v2

@@ IV-A3 Experimental Design
Environmental Parallelization Capability The efficiency of online learning is often dictated by the degree of environment parallelization, which can be limited by the availability of computational resources or the complexity of the dynamics itself, such as large-scale diffusion architectures [53, 100, 52, 99] or the complexity of exteroceptive observations (e.g., high-resolution RGB images).
@@
PPO [77] DDPG [50] TD3 [33] SAC [37] DIPO [103] QVPO [21] DACER [95] GenPO [22]
@@
IV-B Experiment 1: General Performance w.r.t. Various Tasks
@@
In environments with extensive agent parallelization, proximity-based on-policy algorithms, specifically PPO and GenPO, demonstrate superior overall performance. While PPO serves as a robust and high-performance baseline with a Mean Rank (MR) of 1.83, GenPO achieves the highest overall performance ceiling (MR=1.67), securing the top rank in 7 out of 12 tasks (compared to 4 for PPO). Although PPO shows slightly better results in specific tasks like Cartpole, Humanoid, and Cassie, GenPO demonstrates superior scalability and robustness in more complex scenarios. Notably, GenPO achieves substantial performance gains in high-dimensional and challenging environments, particularly in complex humanoid locomotion tasks involving Unitree G1 and H1 robots on rough terrain. These results underscore the potential of on-policy DPRL to transcend the representational limitations of conventional Gaussian policies in complex robotic tasks.
Conversely, off-policy algorithms generally exhibit slower convergence and suboptimal asymptotic performance compared to their on-policy counterparts. This performance gap is primarily attributed to the challenge of tracking rapidly shifting data distributions within replay buffers under high-throughput parallelization [22]. However, we observe a notable advantage of incorporating diffusion policies in complex locomotion tasks. Among the off-policy DPRL methods, DIPO achieved a MR of 4.08, significantly outperforming traditional off-policy algorithms such as DDPG (MR=5.83), TD3 (MR=5.08), and SAC (MR=5.08), which is widely considered a strong classical baseline in off-policy RL. These results demonstrate that the robust expressive capacity of Diffusion Models can effectively address the limitations of conventional off-policy RL, offering a more resilient framework for complex control tasks.
@@
A closer examination of the four DPRL paradigms reveals distinct limitations: DIPO consistently outperforms QVPO and DACER. The suboptimal performance of QVPO (MR=5.83) is likely due to its sensitivity to reward scaling; since the $Q$-related reweighting function is directly coupled with the magnitude of manually designed rewards, it struggles with inconsistent reward landscapes across different tasks. DACER recorded the lowest performance (MR=6.58), failing to converge in the majority of tasks. This failure can be attributed to the inherent high variance of policy gradients [35] and the extreme difficulty of optimizing diffusion models directly through BPTT using noisy Q-values as a supervisory signal, which often leads to unstable gradient dynamics [9, 12].
IV-C Experiment 2: Parallelization Capability
These findings suggest a clear trade-off: while GenPO and PPO are optimized for high-throughput scalability, DIPO is uniquely suited for scenarios where extensive parallelization is computationally prohibitive. This is particularly relevant for training overparameterized models, such as Vision-Language-Action (VLA) architectures [46], where the memory and computational costs of maintaining thousands of concurrent environments are often infeasible.
Furthermore, our analysis highlights a critical gap in current online DPRL literature, which predominantly reports performance at a single, fixed scale [71, 103, 21, 57, 95]. Our results underscore that environment scale is not merely a hyperparameter but a decisive factor in algorithmic ranking. Consequently, future benchmarking in the DPRL domain should incorporate diverse parallelization scales to ensure a comprehensive assessment of both scalability and robustness.
@@ IV-C Experiment 2: Parallelization Capability
DIPO [103] QVPO [21] DACER [95] GenPO [22]
@@
PPO [77] 874.32 1.201 0.555 470.40 1.031 0.517 4.92 1.630 1.287
DDPG [50] 629.96 1.580 1.806 539.40 1.607 2.026 96.04 1.619 1.599
TD3 [33] 580.63 1.595 1.859 559.83 1.635 2.002 90.61 1.652 1.498
SAC [37] 610.05 1.552 1.787 583.63 1.504 1.960 99.49 1.649 1.645
DIPO [103] 546.78 1.518 1.816 530.75 1.641 2.016 97.49 1.548 1.617
QVPO [21] 564.87 1.597 1.820 564.27 1.545 2.045 98.62 1.610 1.581
DACER [95] 644.37 1.569 1.825 552.73 1.587 2.023 88.93 1.530 1.562
GenPO [22] 961.39 0.495 0.344 862.80 0.459 0.400 58.30 0.786 0.547
@@
PPO [77] 423.68 1.377 0.776 789.82 1.330 0.600 858.58 1.357 0.640
DDPG [50] 64.06 1.627 1.996 690.03 1.534 1.814 651.14 1.478 1.821
TD3 [33] 70.54 1.670 1.880 731.15 1.522 1.797 582.41 1.543 1.770
SAC [37] 89.21 1.571 1.860 671.47 1.622 1.815 518.65 1.511 1.869
DIPO [103] 83.36 1.719 1.917 685.57 1.599 1.788 474.08 1.604 1.845
QVPO [21] 86.36 1.580 1.937 704.72 1.568 1.828 510.62 1.490 1.855
DACER [95] 58.93 1.660 2.000 719.18 1.525 1.799 625.11 1.567 1.798
GenPO [22] 869.12 0.596 0.431 568.34 0.850 0.568 493.95 0.992 0.521
@@
IV-D Experiment 3: Diffusion Step Scalability
@@
Conversely, DACER and GenPO, which perform backpropagation through all timesteps during training, exhibit a distinct inverse scalability pattern. While GenPO achieves state-of-the-art performance at lower step counts (KK==5, 10), its performance deteriorates significantly as $K$ increases, eventually suffering a catastrophic drop at KK==50. We attribute this failure to the numerical instability and gradient vanishing inherent in backpropagating through an excessively long denoising chain [9, 70]. Indeed, training diffusion models via BPTT is notoriously difficult, as the recursive nature of the denoising process compounds gradient-related issues as the chain length grows [12]. Interestingly, DACER exhibits the most erratic behavior; despite being a BPTT-based method, it fails to achieve meaningful reward levels regardless of the $K$ value. This suggests that its BPTT objective may fundamentally have a scalability issue with end-to-end diffusion optimization in online DPRL tasks.
IV-E Experiment 4: Cross-embodiment Generalizability
@@
IV-F Experiment 5: OOD Environment Generalizability
@@ V Discussion
V-A Topic 1. What are the current challenges of the online DPRL?
Our empirical analysis in Sec. IV reveals several fundamental bottlenecks inherent in existing online DPRL paradigms. First, DIPO, one of the representative algorithms of action gradient methods, often suffers from suboptimal performance due to the difficulty of obtaining accurate gradients from a potentially biased Q-critic [73, 21]. Specifically, updating actions via deterministic gradients and re-storing them into the replay buffer in DIPO may violate the Markov Decision Process (MDP) assumptions, as the stored transitions no longer reflect the current environment-agent interaction distribution [49].
Second, QVPO, categorized by $Q$-weighting methods, exhibits vulnerability to reward scale sensitivities, as the $Q$-related reweighting function is directly coupled with the magnitude of manually designed rewards. For example, the derived Q-weighted variational objective for diffusion policy training cannot handle negative rewards properly [57].
Third, while the Proximity-based method, GenPO, benefits from large-scale parallelization in simulation, its performance degrades significantly in resource-constrained scenarios (refer to Sec. IV-C). Moreover, the intractability of the exact log-likelihood in diffusion models necessitates crude approximations or specialized architectures like invertible diffusion [94], which imposes a bottleneck on both representational flexibility and inference speed.
Lastly, as demonstrated in Sec. IV-D, BPTT-based methods such as GenPO and DACER encounter severe scalability issues when the number of diffusion steps is increased. This is primarily due to gradient instability—including vanishing and exploding gradient—which inherently arises when backpropagating through an extensively long denoising chain [9, 70]
In summary, while each online DPRL paradigm offers unique theoretical advantages, they collectively face significant hurdles—ranging from gradient inaccuracy and reward-scale sensitivity to computational intractability and backpropagation-related scalability limits. Addressing these intertwined challenges is essential for transitioning from specialized simulation environments to reliable, high-performance robotic control in diverse real-world scenarios. Moreover, as simulation environments become increasingly efficient at supporting large-scale parallelization [62, 31], the demand for robust on-policy algorithms is expected to grow. Given the current scarcity of literature on on-policy DPRL compared to its off-policy counterparts, we encourage future researchers to explore this relatively uncharted territory to unlock the full potential of high-throughput robotic learning.
V-B Topic 2. How do we select the online DPRL algorithms in specific robotic scenarios?
@@
V-C Topic 3. What is the Future Avenue of Online DPRL?
@@
While several studies [73, 60, 43, 74] have explored integrating online DPRL with offline-to-online and model-based approaches, the field remains in its early stages with numerous open challenges. To advance online DPRL toward more general robotic applications, we identify several promising research directions that leverage established RL paradigms, offering guidance for future investigation.
V-C1 Action Chunking and Trajectory Planning
@@
While existing online DPRL methods focus on single-action per-predictions (as policies), real-world robotics often employs planning-based approaches that generate action chunks [107] or multi-step trajectories [46] (as planners). Recent work on Q-Chunking [48] has demonstrated Q-learning extensions to action sequences for offline-to-online RL. Integrating temporal action chunking into online model-free DPRL could bridge reactive control and deliberative planning by leveraging diffusion models’ multi-step prediction capabilities to generate temporally consistent action sequences aligned with RL objectives. This evolution from state-to-action policies to state-to-trajectory planners would enhance sample efficiency and robustness in long-horizon tasks.
V-C2 Safe Reinforcement Learning Integration
@@
Our experimental analysis (Sec IV-F) shows that simulation-trained policies can exhibit overly aggressive behaviors from overfitting to source domains when maximizing rewards without safety constraints. Safe RL approaches—including constraint-based methods (e.g., Constrained Policy Optimization [3]), barrier functions [5], and robust control [65]—could naturally integrate with online DPRL to provide formal safety guarantees during exploration while maintaining diffusion-based action expressiveness, facilitating safer deployment in safety-critical applications.
V-C3 Multi-Agent Reinforcement Learning
@@
Multi-agent scenarios present exponentially greater complexity than single-agent settings due to non-stationarity, credit assignment, and coordination requirements. However, advances in large-scale physics simulators capable of efficiently simulating thousands of agents make online Multi-Agent DPRL (MADPRL) particularly promising. Initial approaches could integrate established MARL algorithms [104, 55, 72] with online DPRL frameworks, while sophisticated extensions might leverage diffusion models’ ability to capture complex joint action distributions for coordination strategies or employ centralized training with decentralized execution paradigms.
V-C4 Inverse Reinforcement Learning from Demonstrations
@@
Inverse Reinforcement Learning (IRL) [69, 108, 39] infers reward functions from expert demonstrations, solving the inverse problem of standard RL. Since most diffusion policies [19, 15, 44] research relies on imitation learning from demonstrations, a natural extension reformulates these as reward learning problems rather than direct behavioral cloning. This “demonstration-to-reward-to-policy” paradigm offers key advantages: learned rewards generalize beyond demonstrated state-action pairs, provide clear objectives for online fine-tuning, and extract value from suboptimal demonstrations. Integrating IRL with online DPRL enables a refined learning approach where offline behavioral priors are optimized through interaction, potentially surpassing both standard imitation and pure offline-to-online approaches.
V-C5 Hierarchical Reinforcement Learning
@@
Hierarchical RL (HRL) [86, 67, 56] decomposes complex decisions into multi-level structures where higher-level policies select among lower-level sub-policies to accomplish long-horizon tasks. This offers compelling synergy with diffusion policies: while diffusion models excel at generating diverse, multimodal actions, their stochastic nature introduces substantial variance. A hierarchical framework could employ a simpler higher-level policy to select among trajectory samples from lower-level diffusion sub-policies, mirroring batch action sampling [97, 21] but within a principled HRL structure. This transforms the exploration-exploitation trade-off from action space to a more manageable sub-policy selection problem, enabling sample-efficient learning of compositional behaviors while preserving diffusion models’ expressive capacity.
VI Conclusion